# Progressive Distillation Based on Masked Generation Feature Method for Knowledge Graph Completion

### v1

@@ Introduction
Knowledge graphs (KGs) are graph-structured knowledge bases, typically composed of triples $(\mathrm{head}\>\mathrm{entity},\>\mathrm{relation},\>\mathrm{tail}\>\mathrm{entity})$, abbreviate as $(h,\>r,\>t)$. Well known are YAGO (Suchanek et al., 2007), Freebase (Bollacker et al., 2008), Wikidata (Vrandečić and Krötzsch, 2014) etc. KGs have proved to be useful in various downstream tasks such as intelligent question answering (Jia et al., 2021; Saxena et al., 2020), recommendation systems (Wang et al., 2019; Cao et al., 2019), semantic search (Xiong et al., 2017; Berant and Liang, 2014) etc. Despite the significant advances that KGs have made for various applications, they still suffer from the problem of incompleteness as the information in the real world continues to grow. Therefore, for the automatic construction of KGs, knowledge graph completion techniques are crucial.
@@
Existing knowledge graph completion (KGC) tasks can generally be divided into two categories: structure-based and description-based methods. Structure-based methods use the topology and triple structure information of the knowledge graph to represent feature vectors of entity relationships, including TransE(Bordes et al., 2013), ConvE(Dettmers et al., 2018), and R-GCN(Schlichtkrull et al., 2018). Description-based methods, use pre-trained language models and introduce semantic descriptions of entities and relations to learn representations, such as commonly used models like KG-BERT (Yao et al., 2017), PKGC (Lv et al., 2022), and LP-BERT (Li et al., 2022). It is evident that with the rise of pre-trained language models (PLM), description-based methods have gradually taken the lead. By utilizing entity and relation semantic descriptions as auxiliary information and deeply mining the potential knowledge in PLM, description-based methods solve the problem of inductive KGC tasks that structure-based methods cannot handle, while achieving significant improvements in transductive KGC tasks. However, while description-based approaches improve performance, they also bring with them the problems of large model parameter numbers and high computational costs, limiting their application in downstream tasks such as real-time recommendation systems etc. Therefore, model lightweighting is essential.
@@
Knowledge distillation (Hinton et al., 2015), which involves the transfer of latent knowledge from a large teacher model to a small student model using soft labels, is a common method for model compression. It has been widely applied in the fields of computer vision (Zhao et al., 2022) and speech recognition (Kurata and Audhkhasi, 2018). In the field of KGC, there are also related research works (Zhu et al., 2022; Wang et al., 2021b ) that employ ensemble models consisting of multiple structure-based KGC models as multi-teacher models to transfer knowledge to student models in order to reduce embedding dimensions. However, to our knowledge, description-based KGC method does not have a corresponding knowledge distillation method, nor can structure-based KGC distillation strategy be directly applied to the description-based KGC method, because intuitively the model architectures of the description-based KGC method and the structure-based KGC method are too different to be directly migrated. Therefore, we believe that the description-based method needs a simple and efficient knowledge distillation framework to fill this gap.
@@
In recent years, KGC method has developed rapidly. The key idea is to map entities and relations in KGs to a continuous vector space as an embedding representation. Among them, structure-based methods focus on representing the feature vectors of a triple through the structural information of the triple itself or the topology of the KGs. For example, TransE (Bordes et al., 2013) models the triple as a relational translation in Euclidean space; Complex (Trouillon et al., 2016) embeds entities and relations in a complex space to deal with asymmetric relations, while RotatE (Sun et al., 2018) models the triple as a relational rotation in a complex space. With the development of deep learning, CNN-based and GNN-based approaches have been proposed in the industry, such as ConvE (Dettmers et al., 2018) and ConvKB (Dai Quoc Nguyen et al., 2018), which use CNN to capture the local structural information of each triple; SACN (Shang et al., 2019) and CompGCN (Vashishth et al., 2019) extract topological information in KGs to represent the triple. With the rise of PLM (BERT (Kenton and Toutanova, 2019), GPT (Radford et al., 2018), etc.), a large number of PLM-based KGC methods have emerged, such as KG-BERT (Yao et al., 2017), StAR (Wang et al., 2021a ), etc. To represent entity and relational embeddings, they introduce natural language descriptions of entities and relations as auxiliary information to mine the potential entity-relational knowledge in pre-trained language models (Petroni et al., 2019).
@@
Knowledge Distillation (KD) (Hinton et al., 2015) is one of the most common techniques in model compression and is widely used in computer vision (Zhao et al., 2022; Yang et al., 2022) and natural language processing (Sun et al., 2019; Sanh et al., 2019; Wang et al., 2023). The core idea of KD is to use the soft label of a large model (teacher model) to guide the learning of a small model (student model). This has the advantage of reducing the computational and storage resource consumption of the model while ensuring performance, thus making the model lighter. In addition to the role of model compression, KD can also improve model performance, and recent studies (Abnar et al., 2020; Kuncoro et al., 2019) have found that KD can transfer inductive biases between neural networks.The self-distillation strategy (Pham et al., 2022) stands out for its effective knowledge transfer approach, enhancing performance through distillation within its own network. Compared with the traditional logits distillation, masked generation feature distillation method we propose can ensure that the teacher model can transfer knowledge more efficiently, so that students can learn more enriched teacher knowledge. At the same time, combined with the progressive distillation strategy, it can ensure that the model parameters is significantly reduced while maintaining the model performance as much as possible.
@@
KGs is a directed relational graph consisting of entities and relations. It can be defined as $\mathcal{G}=\{\mathcal{E},\>\mathcal{R},\>\mathcal{T}\}$, where $\mathcal{E}$ is the set of entities, $\mathcal{R}$ is the set of relations and $\mathcal{T}$ is the set of triples, denoted as $T=\{(h,\>r,\>t)\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$.The link prediction task aims to complete missing triplets based on the existing KGs information. Specifically, under the widely adopted entity ranking evaluation protocol, tail entity prediction infers the tail entity given a head entity and a relation, head entity prediction is similar. In this paper, inverse triplets are set up for each triplet (Dettmers et al., 2018), so only tail entity prediction needs to be performed in the experiments.
@@
BERT (Zhang and Hashimoto, 2021) model leverages the technique of masked language modeling to acquire valuable inductive biases. Inspired by this idea, we propose a masked generation feature distillation (MGFD), speculating that the concept of masking can also facilitate the transfer of more inductive biases from the teacher model in knowledge distillation. When triplets and their text description pass through deep networks, they acquire higher-order semantic information. In traditional feature distillation, the semantic information only consists of the global semantic information of neighboring tokens and the semantic information of the current token (i.e.input token set). In contrast, our proposed MGFD not only incorporates the semantic information within the input token set but also utilizes the inferential operations of deep networks to obtain semantic information from the inferred token set. This enrichment leads to a more comprehensive representation of the generated features, enabling the student model to learn more abundant representations. Consequently, this approach addresses the problem of inefficient knowledge transmission from the teacher model in the context of KGC tasks.
@@
In this section, we present the progressive distillation framework. Specifically, inspired by the idea of layer-by-layer distillation to transfer knowledge from deep models to shallow models (Xue et al., 2023).By gradually reducing the mask ratio and model parameters, it ensures that teacher model knowledge is effectively transferred to student model. This solves the issue of mismatched masking-generated feature information from the teacher model and the expressive capacity of the student model.
@@
TransE(Bordes et al., 2013) - 24.3 4.3 44.1 53.2 27.9 19.8 37.6 44.1
RotatE(Sun et al., 2018) - 47.6 42.8 49.2 57.1 33.8 24.1 37.5 53.3
ConvE(Dettmers et al., 2018) - 43.0 40.0 44.0 52.0 32.5 23.7 35.6 50.1
CompGCN(Vashishth et al., 2019) - 47.9 44.3 49.4 54.6 35.5 26.4 39.0 53.5
@@
KG-BERT(Yao et al., 2017) 110M 21.6 4.1 30.2 52.4 - - - 42.0
MTL-KGC(Kim et al., 2020) 110M 33.1 20.3 38.3 59.7 26.7 17.2 29.8 45.8
C-LMKE(Wang et al., 2022b ) 110M 61.9 52.3 67.1 78.9 30.6 21.8 33.1 48.4
KGLM(Youn and Tagkopoulos, 2022) 355M 46.7 33.0 53.8 74.1 28.9 20.0 31.4 46.8
LP-BERT(Li et al., 2022) 355M 48.2 34.3 56.3 75.2 31.0 22.3 33.6 49.0
StAR(Wang et al., 2021a ) 355M 40.1 24.3 49.1 70.9 29.6 20.5 32.2 48.2
$\mathrm{Baseline}$(Wang et al., 2022a ) 210M 67.1 58.5 73.1 81.7 33.3 24.6 36.2 51.0
@@
Datasets. We experiment on two common KGC benchmark datasets WN18RR(Dettmers et al., 2018) and FB15k-237(Toutanova and Chen, 2015). WN18RR is a subset of WordNet(Miller, 1995), while FB15K-237 is a subset of Freebase. WN18RR and FB15K-237 resolve the test set leakage problem in WN18 and FB15K by eliminating inverse relations. The statistical data is shown in Table 1.
@@
Table 3: The comparison experiment of the distillation strategy on the WN18RR dataset, ”LKD(Hinton et al., 2015)” is the logits distillation, and ”PKD(Sun et al., 2019)” is the teacher model’s middle feature layer, and the feature distillation is performed by layer skipping. “L” is Layer.
@@
Baselines. We select a representative SimKGC (Wang et al., 2022a ) model as the baseline for our distillation framework. The effectiveness of our framework migration is demonstrated by achieving performance breakthroughs on hard-to-breakthrough high metric models.
@@
We reuse the data results from StAR(Wang et al., 2021a ) regarding TransE and obtained the experimental data for other models from their respective papers’ best results. In Table 2, on the WN18RR dataset, the $\mathrm{PMD_{12}}$ during the pre-distillation stage achieve a significant improvement in all metrics without increasing the model parameters, reaching the SOTA level. We attribute this to the MGFD module, which allows the student model to learn rich representation information, thereby enhancing the model’s robustness and significantly improving its overall performance.

### v2

@@ Introduction
Knowledge graphs (KGs) are graph-structured knowledge bases, typically composed of triples $(\mathrm{head}\>\mathrm{entity},\>\mathrm{relation},\>\mathrm{tail}\>\mathrm{entity})$, abbreviate as $(h,\>r,\>t)$. Well known are YAGO (Suchanek, Kasneci, and Weikum 2007), Freebase (Bollacker et al. 2008), Wikidata (Vrandečić and Krötzsch 2014) etc. KGs have proved to be useful in various downstream tasks such as intelligent question answering (Jia et al. 2021; Saxena, Tripathi, and Talukdar 2020), recommendation systems (Wang et al. 2019; Cao et al. 2019), semantic search (Xiong, Power, and Callan 2017; Berant and Liang 2014) etc. Despite the significant advances that KGs have made for various applications, they still suffer from the problem of incompleteness as the information in the real world continues to grow. Therefore, for the automatic construction of KGs, knowledge graph completion techniques are crucial.
@@
Existing knowledge graph completion (KGC) tasks can generally be divided into two categories: structure-based and description-based methods. Structure-based methods use the topology and triple structure information of the knowledge graph to represent feature vectors of entity relationships, including TransE(Bordes et al. 2013), ConvE(Dettmers et al. 2018), and R-GCN(Schlichtkrull et al. 2018). Description-based methods, use pre-trained language models and introduce semantic descriptions of entities and relations to learn representations, such as commonly used models like KG-BERT (Yao, Mao, and Luo 2017), PKGC (Lv et al. 2022), and LP-BERT (Li et al. 2022). It is evident that with the rise of pre-trained language models (PLM), description-based methods have gradually taken the lead. By utilizing entity and relation semantic descriptions as auxiliary information and deeply mining the potential knowledge in PLM, description-based methods solve the problem of inductive KGC tasks that structure-based methods cannot handle, while achieving significant improvements in transductive KGC tasks. However, while description-based approaches improve performance, they also bring with them the problems of large model parameter numbers and high computational costs, limiting their application in downstream tasks such as real-time recommendation systems etc. Therefore, model lightweighting is essential.
@@
Knowledge distillation (Hinton, Vinyals, and Dean 2015), which involves the transfer of latent knowledge from a large teacher model to a small student model using soft labels, is a common method for model compression. It has been widely applied in the fields of computer vision (Zhao et al. 2022) and speech recognition (Kurata and Audhkhasi 2018). In the field of KGC, there are also related research works (Zhu et al. 2022; Wang et al. 2021b) that employ ensemble models consisting of multiple structure-based KGC models as multi-teacher models to transfer knowledge to student models in order to reduce embedding dimensions. However, to our knowledge, description-based KGC method does not have a corresponding knowledge distillation method, nor can structure-based KGC distillation strategy be directly applied to the description-based KGC method, because intuitively the model architectures of the description-based KGC method and the structure-based KGC method are too different to be directly migrated. Therefore, we believe that the description-based method needs a simple and efficient knowledge distillation framework to fill this gap.
@@
In recent years, KGC method has developed rapidly. The key idea is to map entities and relations in KGs to a continuous vector space as an embedding representation. Among them, structure-based methods focus on representing the feature vectors of a triple through the structural information of the triple itself or the topology of the KGs. For example, TransE (Bordes et al. 2013) models the triple as a relational translation in Euclidean space; Complex (Trouillon et al. 2016) embeds entities and relations in a complex space to deal with asymmetric relations, while RotatE (Sun et al. 2018) models the triple as a relational rotation in a complex space. With the development of deep learning, CNN-based and GNN-based approaches have been proposed in the industry, such as ConvE (Dettmers et al. 2018) and ConvKB (Dai Quoc Nguyen, Nguyen, and Phung 2018), which use CNN to capture the local structural information of each triple; SACN (Shang et al. 2019) and CompGCN (Vashishth et al. 2019) extract topological information in KGs to represent the triple. With the rise of PLM (BERT (Kenton and Toutanova 2019), GPT (Radford et al. 2018), etc.), a large number of PLM-based KGC methods have emerged, such as KG-BERT (Yao, Mao, and Luo 2017), StAR (Wang et al. 2021a), etc. To represent entity and relational embeddings, they introduce natural language descriptions of entities and relations as auxiliary information to mine the potential entity-relational knowledge in pre-trained language models (Petroni et al. 2019).
@@
Knowledge Distillation (KD) (Hinton, Vinyals, and Dean 2015) is one of the most common techniques in model compression and is widely used in computer vision (Zhao et al. 2022; Yang et al. 2022) and natural language processing (Sun et al. 2019; Sanh et al. 2019; Wang et al. 2023). The core idea of KD is to use the soft label of a large model (teacher model) to guide the learning of a small model (student model). This has the advantage of reducing the computational and storage resource consumption of the model while ensuring performance, thus making the model lighter. In addition to the role of model compression, KD can also improve model performance, and recent studies (Abnar, Dehghani, and Zuidema 2020; Kuncoro et al. 2019) have found that KD can transfer inductive biases between neural networks.The self-distillation strategy (Pham et al. 2022) stands out for its effective knowledge transfer approach, enhancing performance through distillation within its own network. Compared with the traditional logits distillation, masked generation feature distillation method we propose can ensure that the teacher model can transfer knowledge more efficiently, so that students can learn more enriched teacher knowledge. At the same time, combined with the progressive distillation strategy, it can ensure that the model parameters is significantly reduced while maintaining the model performance as much as possible.
@@
KGs is a directed relational graph consisting of entities and relations. It can be defined as $\mathcal{G}=\{\mathcal{E},\>\mathcal{R},\>\mathcal{T}\}$, where $\mathcal{E}$ is the set of entities, $\mathcal{R}$ is the set of relations and $\mathcal{T}$ is the set of triples, denoted as $T=\{(h,\>r,\>t)\subseteq\mathcal{E}\times\mathcal{R}\times\mathcal{E}\}$.The link prediction task aims to complete missing triplets based on the existing KGs information. Specifically, under the widely adopted entity ranking evaluation protocol, tail entity prediction infers the tail entity given a head entity and a relation, head entity prediction is similar. In this paper, inverse triplets are set up for each triplet (Dettmers et al. 2018), so only tail entity prediction needs to be performed in the experiments.
@@
BERT (Zhang and Hashimoto 2021) model leverages the technique of masked language modeling to acquire valuable inductive biases. Inspired by this idea, we propose a masked generation feature distillation (MGFD), speculating that the concept of masking can also facilitate the transfer of more inductive biases from the teacher model in knowledge distillation. When triplets and their text description pass through deep networks, they acquire higher-order semantic information. In traditional feature distillation, the semantic information only consists of the global semantic information of neighboring tokens and the semantic information of the current token (i.e.input token set). In contrast, our proposed MGFD not only incorporates the semantic information within the input token set but also utilizes the inferential operations of deep networks to obtain semantic information from the inferred token set. This enrichment leads to a more comprehensive representation of the generated features, enabling the student model to learn more abundant representations. Consequently, this approach addresses the problem of inefficient knowledge transmission from the teacher model in the context of KGC tasks.
@@
In this section, we present the progressive distillation framework. Specifically, inspired by the idea of layer-by-layer distillation to transfer knowledge from deep models to shallow models (Xue et al. 2023).By gradually reducing the mask ratio and model parameters, it ensures that teacher model knowledge is effectively transferred to student model. This solves the issue of mismatched masking-generated feature information from the teacher model and the expressive capacity of the student model.
@@
TransE(Bordes et al. 2013) - 24.3 4.3 44.1 53.2 27.9 19.8 37.6 44.1
RotatE(Sun et al. 2018) - 47.6 42.8 49.2 57.1 33.8 24.1 37.5 53.3
ConvE(Dettmers et al. 2018) - 43.0 40.0 44.0 52.0 32.5 23.7 35.6 50.1
CompGCN(Vashishth et al. 2019) - 47.9 44.3 49.4 54.6 35.5 26.4 39.0 53.5
@@
KG-BERT(Yao, Mao, and Luo 2017) 110M 21.6 4.1 30.2 52.4 - - - 42.0
MTL-KGC(Kim et al. 2020) 110M 33.1 20.3 38.3 59.7 26.7 17.2 29.8 45.8
C-LMKE(Wang et al. 2022b) 110M 61.9 52.3 67.1 78.9 30.6 21.8 33.1 48.4
KGLM(Youn and Tagkopoulos 2022) 355M 46.7 33.0 53.8 74.1 28.9 20.0 31.4 46.8
LP-BERT(Li et al. 2022) 355M 48.2 34.3 56.3 75.2 31.0 22.3 33.6 49.0
StAR(Wang et al. 2021a) 355M 40.1 24.3 49.1 70.9 29.6 20.5 32.2 48.2
$\mathrm{Baseline}$(Wang et al. 2022a) 210M 67.1 58.5 73.1 81.7 33.3 24.6 36.2 51.0
@@
Datasets. We experiment on two common KGC benchmark datasets WN18RR(Dettmers et al. 2018) and FB15k-237(Toutanova and Chen 2015). WN18RR is a subset of WordNet(Miller 1995), while FB15K-237 is a subset of Freebase. WN18RR and FB15K-237 resolve the test set leakage problem in WN18 and FB15K by eliminating inverse relations. The statistical data is shown in Table 1.
@@
Table 3: The comparison experiment of the distillation strategy on the WN18RR dataset, ”LKD(Hinton, Vinyals, and Dean 2015)” is the logits distillation, and ”PKD(Sun et al. 2019)” is the teacher model’s middle feature layer, and the feature distillation is performed by layer skipping. “L” is Layer.
@@
Baselines. We select a representative SimKGC (Wang et al. 2022a) model as the baseline for our distillation framework. The effectiveness of our framework migration is demonstrated by achieving performance breakthroughs on hard-to-breakthrough high metric models.
@@
We reuse the data results from StAR(Wang et al. 2021a) regarding TransE and obtained the experimental data for other models from their respective papers’ best results. In Table 2, on the WN18RR dataset, the $\mathrm{PMD_{12}}$ during the pre-distillation stage achieve a significant improvement in all metrics without increasing the model parameters, reaching the SOTA level. We attribute this to the MGFD module, which allows the student model to learn rich representation information, thereby enhancing the model’s robustness and significantly improving its overall performance.